A car moving with a velocity of  10m\s and with a uniform acceleration gains a velocity of 30 m\s after travelling a distance of 4km. The time taken is
 

Grade:7

1 Answers

Arun
5 years ago
Dear student
 
when an object travels in a straight line-

speed= velocity

avg speed=avg velocity
avg velocity = u+v/2
10+30/2
=20m/s
v=s/t
v=20m/s
s=4km(4000m)
t=v/s
=1/200